# Data Quality Issues in Flare Prediction using Machine Learning Models

### v1

@@ 2 Flare Response: Data Source, Quality, and Processing
2.1 Sources of Data Products for Flare Forecasting
2.1.1 GOES X-ray Flare
@@
Figure 2: The 1-minute averaged X-Ray flux in log10 scale and the flare events at May 3rd, 2022. The short, vertical lines represent the event time of flares. Not every local maximum in flux will be labeled as one flare event. The flux and the flare events are the Science-Quality data processed by NCEI.
@@
A notable change between the GOES-R (GOES 16-18) and previous GOES data is that the operational (near-real-time) data prior to GOES-16 had scaling factors applied by SWPC to adjust the GOES 8-15 irradiances to match fluxes from GOES-7. Thus, the flare index before GOES-16 was based on the operational irradiances with scaling factors. To get the true irradiances, the scaling factor of 0.7 (for the XRS-B channel) had to be removed (see Section 2.2).
@@
The operational flare data based on the operational 1-minute X-Ray flux is prepared and stored by SWPC. SWPC provides near-real-time and recently observed X-ray flux data, solar and geomagnetic indices and solar event reports created from preliminary reports. Specifically, the SWPC offers Solar Events Lists (SEL) from 1996 to the present, accessible via its public warehouse (see ftp://ftp.swpc.noaa.gov/pub/warehouse/). The SEL provides daily preliminary lists of solar events, such as bright surges on the limb, filament disappearances, optical flares observed in H-alpha, X-ray flares, and others. By selecting the XRA event type, users can access operational X-ray flare event data dating back to 1996 (SEL). We refer to the flare event lists from the SEL as the SWPC-FTP list. The SWPC-FTP list includes flare times, magnitudes, and the assigned NOAA solar active region numbers. Additionally, the warehouse contains the daily SRS (Solar Region Summary) report, which provides detailed information about the solar active regions, including their positions.
@@
Science-Quality XRS data are produced by NOAA’s National Center for Environmental Information (NCEI) and differ from the operational products used at SWPC in that the data have been preprocessed and incorporate retrospective fixes for issues and outages in the operational product. Specifically, NCEI has reprocessed GOES 8-18 operational flare summary data to produce science-quality flare summary data (see Section 2.2). The GOES 1-7 data will be completed in 2026. The Science-Quality flare summary data is based on the 1-minute Science-Quality X-Ray irradiances data. We refer to this flare catalog as the Science-Quality list in this paper. The Science-Quality list does not include the corresponding active region (AR) number for each flare event. However, NCEI provides Science-Quality flare location information for GOES-R (GOES 16-18) data. The flare location on the solar disk is determined based on the measurements from the high flux XRS-B2 quad diode detector and was calibrated and validated with comparisons to flare locations from the Heliophysics Event Registry (goesr_l2_userguide). And this can be used to match to the ARs.
@@
2.1.2 AIA Flares and SolarSoft Latest Events
@@
Besides GOES X-Ray flare, AIA flares are solar flares detected and characterized based on extreme ultraviolet (EUV) observations from the Atmospheric Imaging Assembly (AIA lemen2012atmospheric) aboard the Solar Dynamics Observatory (SDO pesnell2012solar). The detection process uses a peak-detection algorithm on binned AIA images, primarily in the 193 Åwavelength, to identify the start, peak, end time, and approximate location of the flare. Basically, the AIA-flares detection algorithm detects flares as intervals of positive derivative and negative derivative around a local maximum (Martens2012).
@@
SolarSoft Latest Events(Freeland2002) is another catalog that regularly detects and collects solar flare events. The catalog is primarily designed to assign spatial locations to GOES flare events by using the best available X-ray/EUV full-disk data. Since 2010, AIA 94åhas been the primary data source, and currently, GOES-R series SUVI 93/94 Åserves as a backup when AIA data is unavailable. It is independent of specific instruments (personal communications with Dr. Samuel Freeland).
2.1.3 HEK Repository
@@
The HEK is a comprehensive metadata system designed to catalog and provide access to solar events, facilitating efficient data retrieval for researchers (hek). HEK serves as an integrated event repository rather than a direct data source. It aggregates flare information from multiple sources. HEK incorporates AIA flare detections from automated feature-finding tools applied to 94 ÅEUV images. It also ingests external sources, including the SWPC operational GOES X-Ray flare list and the SSW catalog.
Additionally, the SunPy package allows users to access solar information through its HEK module. According to the official SunPy document for accessing the GOES flare list, as well as the open source code, users can retrieve the desired SWPC operational flare events from the HEK repository by setting the appropriate event type and observatory.
@@
However, the GOES flare list from the SunPy-HEK module exhibits several discrepancies compared to the flare list available in the SWPC open warehouse, particularly in event counts and Active Region (AR) number information. Between January 1, 1997, and July 21, 2024, 289 flares (1.14%) appear in the SunPy-HEK list but are absent from the SWPC-FTP list, while 817 flares (3.15%) recorded by SWPC are not found in the SunPy-HEK list. Additionally, 9842 flare events have differing AR numbers between the two lists. Figure 3 shows the cumulative flare intensity since January 1, 2010, for both the SunPy-HEK and SWPC-FTP flare lists. A notable disparity appears after 2022, due to 728 flare events recorded by SunPy-HEK that are missing, and more than 3,000 flares are mislabeled with AR number 0 in the SunPy-HEK list. The HEK operation team suggested that the difference in event counts may be due to the double retrieval of the same flare after SWPC revisions (personal communication). Thus, we do not recommend using the SunPy-HEK GOES flare list, especially the GOES-R data, for research purposes.
@@ 2.2 Comparison of Science Quality and Operational Flare Response
NCEI produces and restores science-quality GOES XRS L1b and L2 datasets in netCDF format. The science-quality data have been reprocessed from the start of the mission to the present date and incorporate retrospective fixes for issues and outages in the operational product. The XRS L1b data contains high-cadence soft X-ray irradiance measurements, and the L2 data consists of higher-order products such as irradiance averages, flare event summaries, and flare location products. Science-quality L2 data products are created from science-quality L1b data. Up to October 2025, science-quality flare summary products are available for GOES-R (GOES 16, 17, 18) and GOES 8-15 satellites, dating from 1995/01/03 to the present. 1-minute flux average product, another commonly used L2 product, has been processed to science-quality for GOES 8-18 satellites. This paper focuses mainly on the flare summary product, which contains the most widely used flare response information in flare prediction tasks.
The science-quality data XRS have been reprocessed with many corrections, including smoothness, accuracy, scaling factor consistency, bandpass corrections, and improvement of the data quality flag (goes8-15_readme). The SWPC scaling factor has the most crucial effect on the flare summary product. Science-quality XRS irradiances are provided in physical units of W/m2 (watts per square meter), while the operational data before GOES-16 had SWPC-adjusted scaling factors that adjusted the GOES 8–15 irradiances to match those of GOES 7. Since the flare summary product is based on the L1b data using the flare detection algorithm, the threshold for each flare class increased by 42%, meaning that an individual solar flare before GOES-R needed to be 42% larger than a flare after GOES-R to be given the same level of magnitude. Thus, fewer flares were covered in the operational data set during GOES 8-15. The science-quality data were corrected so that they did not contain the SWPC scaling factors. Figure 4 shows the flare magnitude inconsistency in the operational data. Figure 5 presents the flare number difference related to the SWPC rescaling factors between the operational data and science-quality data.
Figure 4: The flare peak flux ratio of SWPC-FTP list to the NCEI science-quality list through 2020/01/01 - 2024/7/21. Before December 2019, the SWPC applied a rescaling factor such that the ratio is centered at 0.7, while the centered ratio is back to 1 as GOES-16 became the primary operational satellite. The colored rectangle shows the data availability time range for GOES 13-18, which differs from the serving time as the primary satellite providing operational data. GOES-16 became the primary geostationary satellite for its operations in December 2019.
@@ 2.3 Cross-Comparison of Science-Quality, Operational, and SSW Flare Lists
To further look at the differences of these flare lists, we then compared the Science-Quality, SWPC-FTP, and SSW flare lists. The Science-Quality catalog was obtained through the NCEI website. The SWPC-FTP catalog is directly downloaded from the SWPC open warehouse due to the inconsistency between the FTP archive and SunPy-HEK records. The SSW flares were downloaded from their web archive (Freeland2002). Additionally, we downloaded the Science-Quality flare location production from NCEI and augmented the Science-Quality list using the pinpoint location information.
The period of interest covers from 2010-01-01 to 2024-12-31. NCEI provides the processed science-quality data for each instrument separately, and the data availability coverage of one instrument usually does not align with the time range when the instrument is in operational use. For example, GOES-16 reached geostationary orbit on 2015-11-29 and started to provide flare event data since 2017-02-09, while GOES-16 did not become NOAA’s primary source of solar X-ray flux until 2019-12-09. GOES-16 was replaced by GOES-18 as the primary satellite in December 2024.
During this time range, GOES-13, -14, -15, and -16 had become the primary satellites. We compiled the flare summary data from different instruments to get the Science-Quality flare list. Since the flare location product is only available for the GOES-R series and the flare location information is crucial in predictive model training, we include the whole time range of the GOES-16 since 2017-02-19. Before GOES-16, we compiled the science-quality data from different GOES instruments based on primary use status (goes1-15_readme). We show the time range of the compiled science-quality flare data in Table 3.
Table 3: The Science-Quality column indicates the periods used in this study to compile science-quality flare events from various satellites. In general, there have been two operational GOES satellites for each instrument, a primary and a secondary. Primary satellite data should be used over secondary satellite data whenever it is available. There will be gaps in good data for the primary satellites due to eclipses and in-flight calibrations. Satellites are listed in the table only when there are changes. Data courtesy: Kimberly Moreland
@@
During this time range, there are 33215 Science-Quality records, distributed into 139 X-flares and 2497 M-flares. We also downloaded 27586 SWPC-FTP and 33489 SSW flares. We apply minimum data cleaning to these lists, only ensuring the validity of start time, peak time, and flare magnitude features. Table 4 presents the number of flares with different magnitudes in our compiled Science-Quality, SWPC-FTP, and SSW lists.
@@
For location information, Science-Quality and SSW have pinpoint flare location coordinates, and SWPC-FTP and SSW have NOAA AR numbers information.
We then attempt to match each Science-Quality flare with corresponding events in the SWPC-FTP and SSW catalogs based on temporal characteristics (i.e., flare peak time) and flare intensity. Specifically, we define tolerance thresholds for both peak time and the logarithm (base 10) of the flare intensity. For a given flare, we first filter candidates from the comparison catalog whose absolute difference in peak time falls within the defined time tolerance window. We then further refine the matches by selecting only those flares whose difference in log10 intensity is within the specified threshold. In cases where multiple matches satisfy these criteria, the flare with the closest peak time is selected. Figure 6 shows the fraction of matched flares under various tolerance settings.
@@
One important consideration is that, before the designation of GOES-16 as the primary satellite, flare intensities reported in the SWPC-FTP catalog were systematically affected by the SWPC scaling factor. This results in reported $\log_{10}$ intensities that are approximately 0.15 lower than their true values. The extent to which the SSW catalog inherits this bias is not explicitly documented. However, our analysis shows that the SSW catalog exhibits a level of similarity to SWPC-FTP that exceeds the similarity observed between any other pair of catalogs. To evaluate this, we first matched flares between the Science-Quality and SSW lists using only peak time with a strict tolerance of zero. For these temporally matched events, we computed the $\log_{10}$ intensity differences across two periods: before 2019-12-07 (prior to GOES-16 availability) and after 2019-12-09 (once GOES-16 data were adopted operationally). Figure 7 shows the resulting histograms. Before GOES-16 became operational, the intensity differences clustered around 0.15, consistent with the expected offset introduced by the SWPC scaling factor. After 2019-12-07, the differences are centered near zero, indicating the removal of this bias. Together with the matching-proportion pattern shown in Figure 6, these results demonstrate that the SSW catalog is not free from the SWPC scaling prior to GOES-16 and, more generally, that SSW aligns more closely with SWPC-FTP than with the Science-Quality catalog. Therefore, the SSW list should be regarded and processed as an operational dataset.
Consequently, for events occurring before 2019-12-09, we adjusted the flare log intensity for SWPC-FTP and SSW by removing the SWPC scaling factor in log scale. In this study, we adopt a peak time tolerance of 15 minutes and a log 10 intensity tolerance of 0.3 as the criteria for identifying corresponding flare events across different catalogs. As shown in Figure 6, when the log intensity tolerance is set over 0.3, the fraction of matched flares remains relatively stable. Figure 8 shows the distribution of Science-Quality flares that fail to match an event in either the operational SWPC-FTP or SSW catalogs. Only 58 flares of class M or higher (2 X-flares) are unmatched, and the overall discrepancy decreases substantially after 2019-12-07, when GOES-R satellites became primarily operational.

### v2

@@ 2 Flare Response: Data Source, Quality, and Processing
2.1 Sources of Data Products for Flare Events
2.1.1 GOES X-ray Flare Data
Geostationary Operational Environmental Satellites (GOES) are operated by NOAA (National Oceanic and Atmospheric Administration ), and observe the Sun in X-ray spectrum. GOES flare event data, 1-minute average X-Ray flux data, are one of the most commonly used resources in flare forecasting tasks and is available from multiple organizations. The GOES X-Ray Sensor (XRS) measurements have been a crucial component of space weather operations since 1975. XRS measurements are in two bandpass channels commonly referred to as the XRS-A (0.05-0.4 nm) and XRS-B (0.1-0.8 nm).
A sudden yet persistent increase in flux is detected as a likely flare. For GOES-R (GOES 16-18) data, the detection algorithm uses a variety of criteria, including an exponential fit, to determine that a flare has begun. Specifically, the algorithm collects a small block of 1-minute averaged XRS-B flux data and smooths the data by a small boxcar filter. The algorithm then checks whether the flux shows signs of a significant rise by evaluating two key features: the presence of an inflection point, where the curve transitions to a faster upward slope, and whether the rise exceeds a defined threshold based on the standard deviation of earlier values. If these conditions are met, it fits the smoothed data to an exponential function. The exponential curve must also show a strong, concave upward shape based on the configurable parameters. If all these criteria are satisfied, the algorithm declares a flare start. The flare peak time is recorded when the flux reaches its maximum. The end time of a flare is declared when the X-ray flux has declined to halfway between the flare peak and the background level (Machol et al., ).
@@ 2.1.1 GOES X-ray Flare Data
Figure 2: The 1-minute averaged X-Ray flux in log10 scale and the flare events at May 3rd, 2022. The short, vertical lines represent the event time of flares. Not every local maximum in flux will be labeled as one flare event. The flux and the flare events are the Science-Quality data processed by the NOAA National Centers for Environmental Information (NCEI).
@@
A notable difference between the GOES-R (GOES 16-18) and previous GOES data is that the operational (near-real-time) data prior to GOES-16 had scaling factors applied by the SWPC (Space Weather Prediction Center) to adjust the GOES 8-15 irradiance to match fluxes from GOES-7. Thus, the flare index before GOES-16 was based on the operational irradiance with scaling factors. To obtain the true irradiance, the scaling factor of 0.7 (for the XRS-B channel) have to be removed (see Section 2.2).
@@
Operational flare data based on the 1-minute X-Ray flux is prepared and stored by the SWPC. SWPC provides near-real-time and recently observed X-ray flux data, solar and geomagnetic indices, and solar event reports created from preliminary reports. Specifically, SWPC offers Solar Events Lists (SEL) from 1996 to the present, accessible through its public warehouse (see ftp://ftp.swpc.noaa.gov/pub/warehouse/). The SEL provides daily preliminary lists of solar events, including bright surges on the limb, filament disappearances, H-alpha optical flares, X-ray flares, and others. By selecting the XRA event type, users can access operational X-ray flare event data dating back to 1996 (SWPC, ). We refer to the flare event lists from the SEL as the SWPC-FTP list. The SWPC-FTP list includes flare times, magnitudes, and the assigned NOAA solar active region numbers. Additionally, the warehouse contains the daily SRS (Solar Region Summary) report, which provides detailed information about the solar active regions, including their positions.
@@
Science-Quality XRS data is produced by NOAA’s National Center for Environmental Information (NCEI) and differ from the operational products used at SWPC in that the data have been preprocessed and incorporate retrospective fixes for issues and outages in the operational product. Specifically, NCEI has reprocessed GOES 8-18 operational flare summary data to produce science-quality flare summary data (see Section 2.2). The GOES 1-7 data will be completed in 2026. The Science-Quality flare summary data are based on 1-minute Science-Quality X-ray irradiance data. We refer to this flare catalog as the Science-Quality list in this paper. The Science-Quality list does not include the corresponding active region (AR) number for each flare event. However, NCEI provides Science-Quality flare location information for GOES-R (GOES 16-18) data. The flare location on the solar disk is determined based on the measurements from the high flux XRS-B2 quad diode detector and was calibrated and validated with comparisons to flare locations from the Heliophysics Event Registry (https://www.lmsal.com/heksearch/, Machol et al. ). And this can be used to match flares to the corresponding ARs. Table 2 summarizes the different GOES X-ray flare sources.
@@
2.1.2 AIA Flare Catalog, SolarSoft Latest Events, and HEK Repository
@@
Besides GOES X-Ray flare, AIA flares are solar flares detected and characterized based on extreme ultraviolet (EUV) observations from the AIA (Lemen et al. (2012)) aboard SDO( Pesnell et al. (2012a)). The detection process uses a peak-detection algorithm applied to binned AIA images, primarily at 193 Å, to identify the start, peak, and end times and the approximate location of the flare. The AIA flare detection algorithm detects flares as intervals of positive derivative and negative derivative around a local maximum (Martens et al., 2012).
@@
SolarSoft Latest Events (Freeland, 2002) is another catalog that regularly detects and collects solar flare events. The catalog is primarily designed to assign spatial locations to GOES flare events using the best available full-disk X-ray/EUV data. Since 2010, AIA 94åhas been the primary data source, and currently, GOES-R series SUVI 93/94 Åserves as a backup when AIA data is unavailable. It is independent of specific instruments (based on personal communications with Dr. Samuel Freeland).
@@
The HEK (Heliophysics Events Knowledgebase) is a comprehensive metadata system designed to catalog and provide access to solar events, facilitating efficient data retrieval for researchers (Hurlburt et al., 2012). The HEK serves as an integrated event repository rather than a direct data source. It aggregates flare information from multiple sources. The HEK incorporates AIA flare detections from automated feature-finding tools applied to 94 ÅEUV images. It also ingests external sources, including the SWPC operational GOES X-Ray flare list and the SSW catalog.
@@
The SunPy package enables users to access solar data via its HEK module. According to the official SunPy documentation and the open-source code, users can retrieve the desired SWPC operational flare events from the HEK repository by specifying the appropriate event type and observatory. However, the GOES flare list from the SunPy-HEK module exhibits several discrepancies compared to the flare list available in the SWPC open warehouse, particularly in event counts and Active Region (AR) number information. Between January 1, 1997, and July 21, 2024, 289 flares (1.14%) appear in the SunPy-HEK list but are absent from the SWPC-FTP list, while 817 flares (3.15%) recorded by SWPC are not found in the SunPy-HEK list. Moreover, 9842 flare events have differing AR numbers between the two lists. Figure 3 shows the cumulative flare intensity since January 1, 2010, for both the SunPy-HEK and SWPC-FTP flare lists. A notable disparity appears after 2022, due to 728 flare events recorded by SunPy-HEK that are missing, and more than 3,000 flares are mislabeled with AR number 0 in the SunPy-HEK list. The HEK operation team suggested that the difference in event counts may be due to the double retrieval of the same flare after SWPC revisions (based on personal communications). Thus, we do not recommend using the SunPy-HEK GOES flare list, especially the GOES-R data, for research purposes.
@@ 2.2 Comparison of Science Quality and Operational Flare Response
NCEI produces and restores science-quality GOES XRS L1b and L2 datasets in netCDF format. The science-quality data have been reprocessed from the start of the mission to the present date and incorporate retrospective fixes for issues and outages in the operational product. The XRS L1b data contain high-cadence soft X-ray irradiance measurements, and the L2 data consist of higher-order products such as irradiance averages, flare event summaries, and flare location products. Science-quality L2 data products are created from science-quality L1b data. Up to October 2025, science-quality flare summary products are available for GOES-R (GOES 16, 17, 18) and GOES 8-15 satellites, dating from 1995/01/03 to the present. The 1-minute flux average product, another commonly used L2 product, has been processed to science-quality for GOES 8-18 satellites. This paper focuses primarily on the flare summary product, which contains the most widely used flare-response information for flare-prediction tasks.
The science-quality data XRS have been reprocessed with numerous corrections, including smoothing, accuracy, scaling factor consistency, bandpass corrections, and improvements to the data quality flag (Mothersbaugh III et al., 2023). The SWPC scaling factor has the most crucial effect on the flare summary product. Science-quality XRS irradiances are provided in physical units of W/m2 (watts per square meter), while the operational data before GOES-16 had SWPC-adjusted scaling factors that adjusted GOES 8–15 irradiance to match those of GOES 7. Since the flare summary product is based on the L1b data using the flare detection algorithm, the threshold for each flare class increased by 42%, meaning that an individual solar flare before GOES-R needed to be 42% larger than a flare after GOES-R to be given the same level of magnitude. Thus, fewer flares were covered in the operational data set during GOES 8-15. The science-quality data were corrected to exclude the SWPC scaling factors. Figure 4 shows the flare magnitude inconsistency in the operational data. Figure 5 presents the flare number difference related to the SWPC rescaling factors between the operational data and science-quality data.
Figure 4: The flare peak flux ratio of SWPC-FTP list to the NCEI science-quality list through 2020/01/01 - 2024/7/21. Before December 2019, the SWPC applied a rescaling factor so that the ratio was centered at 0.7, whereas the centered ratio returned to 1 as GOES-16 became the primary operational satellite. The colored rectangle shows the data availability time range for GOES 13-18, which differs from the serving time as the primary satellite providing operational data. GOES-16 became the primary geostationary satellite in December 2019.
@@ 2.3 Cross-Comparison of Science-Quality, Operational, and SSW Flare Lists
To further investigate the differences of these flare lists, we compared the Science-Quality, SWPC-FTP, and SSW flare lists. The Science-Quality catalog was obtained through the NCEI website. The SWPC-FTP catalog is directly downloaded from the SWPC open warehouse due to the inconsistency between the FTP archive and SunPy-HEK records. The SSW flares were downloaded from their web archive (Freeland, 2002). In addition, we downloaded the Science-Quality flare location dataset from NCEI and augmented the Science-Quality list with pinpoint location information. Finally, Science-Quality and SSW have pinpoint flare location coordinates, while SWPC-FTP and SSW have NOAA AR numbers.
The period of interest is from 2010-01-01 to 2024-12-31. NCEI provides processed science-quality data for each instrument separately, and the data availability coverage of one instrument usually does not align with the time range during which the instrument is in operational use. For example, GOES-16 reached geostationary orbit on 2015-11-29 and started to provide flare event data since 2017-02-09, while GOES-16 did not become NOAA’s primary source of solar X-ray flux until 2019-12-09. GOES-16 was replaced by GOES-18 as the primary satellite in December 2024. During this period, GOES-13, -14, -15, and -16 became the primary satellites. We compiled the flare summary data from different instruments to obtain the Science-Quality flare list. Since the flare location product is available only for the GOES-R series and flare location information is crucial for training predictive models, we include the full time range of GOES-16 since 2017-02-19. Before GOES-16, we compiled science-quality data from various GOES instruments based on their primary use status (Machol et al., ). We show the time range of the compiled science-quality flare data in Table 3.
Table 3: The Science-Quality column indicates the periods used in this study to compile science-quality flare events from various satellites. In general, there have been two operational GOES satellites for each instrument, a primary and a secondary. Primary satellite data should be used in preference to secondary satellite data whenever available. There will be gaps in good data for the primary satellites due to eclipses and in-flight calibrations. Satellites are listed in the table only when changes occur. Data courtesy: Kimberly Moreland.
@@
During this time range, there are 33,215 Science-Quality records, distributed into 139 X-flares and 2,497 M-flares. We also downloaded 27,586 SWPC-FTP, and 33,489 SSW flares. We perform minimum data cleaning to these lists, only ensuring the validity of start time, peak time, and flare magnitude features. Table 4 presents the number of flares with different magnitudes in our compiled Science-Quality, SWPC-FTP, and SSW lists.
@@
We then attempt to match each Science-Quality flare with corresponding events in the SWPC-FTP and SSW catalogs based on temporal characteristics (i.e., flare peak time) and flare intensity. Specifically, we define tolerance thresholds for both peak time and the logarithm (base 10) of the flare intensity. For a given flare, we first filter the comparison catalog to exclude candidates whose absolute difference in peak time exceeds the defined time tolerance window. We further refine the matches by selecting only flares whose log10 intensity difference is within the specified threshold. When multiple matches satisfy these criteria, the flare with the closest peak time is selected. Figure 6 shows the fraction of matched flares under various tolerance settings.
@@
One important consideration is that, before the designation of GOES-16 as the primary satellite, flare intensities reported in the SWPC-FTP catalog were systematically affected by the SWPC scaling factor. This results in reported $\log_{10}$ intensities that are approximately 0.15 lower than their true values. The extent to which the SSW catalog inherits this bias is not explicitly documented. However, our analysis shows that the SSW catalog exhibits greater similarity to SWPC-FTP than to any other pair of catalogs. To evaluate this, we first matched flares between the Science-Quality and SSW lists using only peak time with a strict tolerance of zero. For these temporally matched events, we computed the $\log_{10}$ intensity differences across two periods: before 2019-12-07 (prior to GOES-16 availability) and after 2019-12-09 (once GOES-16 data were adopted operationally). Figure 7 shows the resulting histograms. Before GOES-16 became operational, the intensity differences clustered around 0.15, consistent with the expected offset introduced by the SWPC scaling factor. After 2019-12-07, the differences are centered near zero, indicating that this bias has been removed. Together with the matching-proportion pattern shown in Figure 6, these results demonstrate that the SSW catalog is not free from the SWPC scaling prior to GOES-16 and, more generally, that SSW aligns more closely with SWPC-FTP than with the Science-Quality catalog. Therefore, the SSW list should be treated as an operational dataset.
Consequently, for events occurring before 2019-12-09, we adjusted the flare log intensity for SWPC-FTP and SSW by removing the SWPC scaling factor in log scale. In this study, we adopt a 15-minute peak-time tolerance and a 0.3 log10 intensity tolerance as criteria for identifying corresponding flare events across different catalogs. As shown in Figure 6, when the log intensity tolerance is set over 0.3, the fraction of matched flares remains relatively stable. Figure 8 shows the distribution of Science-Quality flares that fail to match an event in either the operational SWPC-FTP or SSW catalogs. Only 58 flares of class M or higher (2 X-flares) are unmatched, and the overall discrepancy decreases substantially after 2019-12-07, when the GOES-R satellites became primarily operational.